OpenStudy (anonymous):

Write the single radical expression for : ^4√x*^5√2x

jimthompson5910 (jim_thompson5910):

\[\large \sqrt[4]{x^5}\sqrt[5]{2x}\] \[\large (x^5)^{\frac{1}{4}}(2x)^{\frac{1}{5}}\] \[\large (x^5)^{\frac{5}{20}}(2x)^{\frac{4}{20}}\] \[\large \left((x^5)^5(2x)^4\right)^{\frac{1}{20}}\] \[\large \sqrt[20]{(x^5)^5(2x)^4}\] \[\large \sqrt[20]{x^{25}(16x^4)}\] \[\large \sqrt[20]{16x^{29}}\] So \[\large \sqrt[4]{x^5}\sqrt[5]{2x}=\sqrt[20]{16x^{29}}\]